OpenStudy (anonymous):

Distribute: 5(2x + 3) 10x + 15 7x + 8 10x + 25 7x + 3

OpenStudy (anonymous):

Write it properly..

OpenStudy (anonymous):

After 10x + 15, what have you written??

OpenStudy (anonymous):

That is how it's written.

OpenStudy (anonymous):

Distribute: 5(2x + 3) A 10x + 15 B 7x + 8 C 10x + 25 D 7x + 3

OpenStudy (anonymous):

What is the actual question? Or can we see the screenshot or pic of that??

OpenStudy (anonymous):

Oh, those are the options.. write it clearly buddy..

OpenStudy (anonymous):

Just distribute 5 to brackets..

OpenStudy (anonymous):

Multiply both sides by 5. \[2x*5\] \[3\times5\]

OpenStudy (anonymous):

\[a(b+c) = ab + ac \quad \quad \text{Distributive Property}\]

OpenStudy (anonymous):

Still in problem??

OpenStudy (anonymous):

There??

OpenStudy (anonymous):

I have to go now, If still you are in trouble, then @ganeshie8 can help you.. :)

OpenStudy (anonymous):

You dont need Ganeshie8 when NightWatcher is here!

OpenStudy (anonymous):

Or @Night-Watcher will definitely help you.. :P

OpenStudy (anonymous):

Ill help!

OpenStudy (anonymous):

I know you will reply that's why I write "can" there.. :P

OpenStudy (anonymous):

@waterineyes Lol

OpenStudy (anonymous):

he he he.. :)

OpenStudy (anonymous):

Are you still having trouble? @CrazyCountryGirl

OpenStudy (anonymous):

May be @CrazyCountryGirl is out of the country for a while.. :)

OpenStudy (anonymous):

@waterineyes Lmao

OpenStudy (anonymous):

I'm here sorry. :)

OpenStudy (anonymous):

Oh, then I think @Night-Watcher has posted the question and I am helping him.. :P

OpenStudy (anonymous):

@CrazyCountryGirl Basically what were trying to say is, do you still need help? Because we threw out ways to solve your question and were not getting much of a response.

OpenStudy (anonymous):

You are here but you are not mentally here..

OpenStudy (anonymous):

I am here, I'm still confused, what do we do with x?

OpenStudy (anonymous):

x will remain x..

OpenStudy (anonymous):

Just a digit will get attached to it.. suppose I am multiplying 3 with a number n: then answer would be : 3n

OpenStudy (anonymous):

If I were adding 3 to the number n then it would be : n + 3

OpenStudy (anonymous):

With that said your answer should be clear in front of you.

OpenStudy (anonymous):

would be a or c?

OpenStudy (anonymous):

Do you know the table for 5 ??

OpenStudy (anonymous):

yes. im not stupid... o_o

OpenStudy (anonymous):

I mean suppose there are 3 students, and I am DISTRIBUTING 5 chocolates to them, then how many total chocolates I am having??

OpenStudy (anonymous):

I asked because I am stupid, and I need your help while using that.. :P

OpenStudy (anonymous):

@CrazyCountryGirl if you want to get rid of your confusion and doubts, then you should interact with us.. :)

OpenStudy (anonymous):

A closed question will mean then you have understood the question properly.. Am I right @CrazyCountryGirl

OpenStudy (anonymous):

*that not then there..

OpenStudy (anonymous):

Can you write today what you are writing?? @Night-Watcher :P

OpenStudy (anonymous):

Yeah >.<...If @CrazyCountryGirl doesn't want to interact...i guess we are done here...

OpenStudy (anonymous):

Thanks but I figured it out, I didn't get how you were explaining. The problem was really simple I read it wrong. :)

OpenStudy (anonymous):

He he he... You are blaming your teacher?? :P

OpenStudy (anonymous):

yep. :)

OpenStudy (anonymous):

\[5(2x+3 ) = 5(2x) + 5(3) = \color{green}{10x + 15}\]

OpenStudy (anonymous):

I have written the formula above, all you have to do is to use that in a proper manner.. :P

OpenStudy (anonymous):

That's what I got. You should delete that though. If another person has this problem then they can get the answer without knowing what to do.

OpenStudy (anonymous):

I bet they can't..

OpenStudy (anonymous):

I have seen people getting afraid from brackets.. () :P

OpenStudy (anonymous):

yeah they could...

OpenStudy (anonymous):

You are 77, you are thinking your way.. :)

OpenStudy (anonymous):

I will definitely tag you at the post where I can prove this.. Check this with your eyeballs..

OpenStudy (anonymous):

@Night-Watcher is also here, you can ask him too, may be I am wrong..

OpenStudy (anonymous):

@waterineyes is certainly correct.